OpenStudy (jennychan12):

Find the area between the curves x = y^3 and x = y^2

OpenStudy (jennychan12):

is it just \[\int\limits_{0}^{1} x^\frac{ 1 }{ 3 } - x^\frac{ 1 }{ 2 } dx\] ?

OpenStudy (anonymous):

So first of all, where do they intersect?

OpenStudy (jennychan12):

oh sorry the limits are from 0 to 1

OpenStudy (anonymous):

Set x equations equal to each other: y³ = y² From there, we can see that y = 0 and 1 Set integration: \(\displaystyle \int_0^1 y^2 - y^3 \space dy\)

OpenStudy (anonymous):

You know, it's just as valid, and much easier, to integrate with respect to \(y\).

OpenStudy (jennychan12):

ohh, ok

OpenStudy (jennychan12):

-_- the area is 0??

OpenStudy (anonymous):

No. How did you get 0, though?

OpenStudy (jennychan12):

oh wait whoops sorry. my mistake.

OpenStudy (jennychan12):

it's 1/12

OpenStudy (jennychan12):

lol

OpenStudy (jennychan12):

wait one question.... sorry, why is it y^2 - y^3 ? i thought it was top curve - bottom curve..?

OpenStudy (anonymous):

Because in between y = 0 and y =1, y² is larger than y³ Let's saying that y = 1/2. y² = 1/4 y³ = 1/8 We can see that y² is larger than y³ in this interval so y² is top curve.

OpenStudy (jennychan12):

oh ok, thanks again :)

OpenStudy (anonymous):

No problem.